# ON SUBGRADIENTS OF CONVEX FUNCTIONS AND ORLICZ PSEUDO-NORMS FOR VECTOR-VALUED FUNCTIONS

### v2

@@ 2. Differentiability and Subgradients of Convex Functions
Hence this set is Borel measurable, and the gradient function $\nabla L(x)$ is Borel measurable as well on $E$. In fact, the function $\nabla L$ is continuous on $E$ (cf. [11]). This implies that if $L$ is differentiable everywhere, then it is $C^{1}$-smooth.
@@
$|y|\,=\,\frac{1}{r}\,\sup_{|h|\leq r}\left<y,h\right>\,\leq\,\frac{1}{r}\,\sup_{|h|\leq 1}(L(x+h)-L(x)),\quad r>0.$ (2.2)
@@ 3. Existence of Measurable Subgradients
This means that the real-valued function $x\rightarrow h_{T(x)}(\theta)$ is upper semi-continuous on $\Omega$ for any $\theta\in{\mathbb{R}}^{n}$, or equivalently, for any $\theta\in S^{n-1}$. Using the identification of ${\mathbb{K}}_{n}$ and $\widetilde{\mathbb{K}}_{n}$, we may consider a map $\widetilde{T}:\Omega\rightarrow\widetilde{\mathbb{K}}_{n}$ and say that it is upper semi-continuous, if, for any $\theta\in S^{n-1}$,
@@
$C=\big\{f\in S^{n-1}:f(\theta_{i})<c_{i}\ \ i=1,\dots,n\big\},\quad\theta_{i}\in S^{n-1},\ c_{i}\in{\mathbb{R}}.$
@@ 4. Remarks on the Construction of Measurable Subgradients
As mentioned in [3] with reference to some recent and old publications, evaluating a subgradient may be a challenging task; this diffculty has motivated the development of numerous subdifferential approximations. In Proposition 3.1 one of those constructions is described using formula (3.1). However, a preferable approach would involve directional derivatives $L(x,\theta)$ defined in (2.3) (which are treated as more accessible quantities). The authors of [3] consider and solve this problem in dimension $n=2$.
@@
Hence, by Jensen’s inequality
@@ 5. The $\Delta_{2}$-Condition
First, let us note that the $\Delta_{2}$-condition is equivalent to the property that the function
@@
$c)$ $R$ is sub-multiplicative: $R(rs)\leq R(r)R(s)$ for all $r,s\geq 0$;
$d)$ $R$ has a sub-polynomial growth: putting $p_{-}=R^{\prime}(1-)$ and $p_{+}=R^{\prime}(1+)$, we have
$R(r)\leq r^{p_{-}}\ (0\leq r\leq 1),\qquad R(r)\leq r^{p_{+}}\ (r\geq 1).$ (5.4)
Proof. The properties a)b)a)-b) are clear. For $c)$, we have, for any $x\neq 0$ and $r,s>0$,
$\frac{L(rsx)}{L(x)}=\frac{L(r(sx))}{L(sx)}\frac{L(sx)}{L(x)}\leq R(r)R(s).$
Taking the supremum over all admissible $x$ on the left-hand side, the claim in $c)$ follows.
This property implies that, for any real $a>0$ and any integer $m\geq 0$,
@@
Hence in the limit (5.6) yields the first bound in (5.4).
Similarly, if $0<r<1$ and $0<a<1$, (5.5) remains to hold and yields (5.6). For $a=1-\varepsilon$ with $\varepsilon\downarrow 0$, by the Taylor expansion
$R(a)=1-R^{\prime}(1-)\,\varepsilon+o(\varepsilon),\quad\log R(a)=-R^{\prime}(1-)\,\varepsilon+o(\varepsilon),$
in the limit (5.6) yields the second bound in (5.4). ∎
Note that the second polynomial bound in (5.4) implies
@@
with $C=\sup_{|x|=1}L(x)$ and $p=p_{+}$.
One can now turn to the characterization of the $\Delta_{2}$-condition. Recall that $L^{\prime}(x,x)$ denotes the directional derivative of $L$ in the direction of $x$ defined in $(2.3)$.
@@
In this case, both suprema are at most $p_{+}=R_{L}^{\prime}(1+)$. If $L$ is differentiable, this condition simplifies to
@@
Let us now see that the finiteness of both suprema in (5.7) is necessary for the $\Delta_{2}$-condition. Assume that (5.1) holds true and put $p=p_{+}$. By the second inequality in (5.4),
@@
Recall that, by the convexity of $L$, for all $h\in{\mathbb{R}}^{n}$ and $y\in\partial L(x)$,
@@ 6. A Two-Sided $\Delta_{2}$-Condition
Again, let $L$ be a non-negative convex function on ${\mathbb{R}}^{n}$ such that $L(0)=0$, $L(x)>0$ for $x\neq 0$. In order to characterize the possible behavior of $L(x)$ for large $x$, Krasnosel’skii and Rutickii also discussed in [4] related properties such as the $\Delta^{\prime}$-condition and the $\Delta_{3}$-condition. Here we focus on a formally different class of convex functions, whose behaviour can be controlled not only at infinity, but also around zero.
@@
It was mentioned in [4] that, if $L$ satisfies (6.1) for all $x$ large enough, it is possible to redefine $L$ near zero in such a way that this inequality would hold for all $x$. The new function is thus equivalent to the original one (in the sense of [4]), and therefore both functions lead to the identical associated Orlicz spaces. Nevertheless, this two-sided $\Delta_{2}$-condition introduces some new remarkable features to the analysis of convex functions.
@@
is finite on the positive half-axis. We call it the Young function associated to $L$. Basic properties of $\Phi_{L}$ are identical to those of $R_{L}$. The next statement is a counterpart of Proposition 5.1 and does not require an additional proof.
@@
As a closely related function, also introduce
@@
Let us record without prroof its basic properties.
@@
Necessarily $p=p_{-}=p_{+}$. Typical examples of such functions are given by $L(x)=\|x\|^{p}$, $x\in{\mathbb{R}}^{n}$, with parameter $p\geq 1$, where $\|\cdot\|$ is an arbitrary norm on ${\mathbb{R}}^{n}$.
@@ 7. Characterization of the Two-Sided $\Delta_{2}$-Conditions
Proof. Suppose that $p=p_{-}>1$ and let $r\geq 1$, $x\in{\mathbb{R}}^{n}$. Using the inequality $L(sx)\leq s^{p}L(x)$ holding with $s\in[0,1]$ as in (6.3), we have
@@
Changing the variable $rx=y$, we get $L^{*}(r^{p-1}y)\leq r^{p}L^{*}(y)$. The replacement $r^{p-1}=s$ yields $L^{*}(sy)\leq s^{q_{+}}L^{*}(y)$, or equivalently
@@
This implies that $L^{*}$ satisfies the two-sided $\Delta_{2}$-condition and implies the first inequality in (7.3).
A similar argument based on (6.3) for the range $0\leq r\leq 1$ leads to
@@
and gives the second inequality in (7.3). ∎
Remark 7.3. If $L(x)\geq c|x|$ for all $x\in{\mathbb{R}}^{n}$ with some constant $c>0$, then $L^{*}(x)=0$ for sufficiently small $|x|$. Hence, in this case we may not speak about the $\Delta_{2}$-condition for the Legendre transform of $L$. For an illustration, let us return to Example 6.5 with $p=1$ and consider the convex function
@@ 8. Luxemburg and Orlicz Pseudo-Norms for Vector-Valued Functions
when the value of $r$ under the infimum sign exists, and otherwise $\|u\|_{L}=\infty$.
@@
Note that $\|u\|_{L}=0$ implies that that $u=0$ with $\lambda$-probability 1. Indeed, suppose that $u\neq 0$ on a set $A\subset\Omega$ of positive $\lambda$-probability. Then, using $L(x)>0$ for all $x\neq 0$, we would have $L(\alpha u)\uparrow\infty$ as $\alpha\uparrow\infty$ on the set $A$, so that $\int L(\alpha u)\,d\lambda\uparrow\infty$, or equivalently $\int L(u/r)\,d\lambda\uparrow\infty$ as $r\downarrow 0$. According to (8.1), we then get $\|u\|_{L}>0$.
@@
for any $r>0$, and hence $\|\widetilde{v}\|_{L^{*}}\leq\|v\|_{L^{*}}$. So, after this modification of $v$ the integrals in (8.2) may only increase, while the norm of the modified $v$ may decrease. That is,
@@
Note that $|u|_{L}=0$ if and only if $\left<u,v\right>\leq 0$ $\lambda$-a.e. whenever $\|v\|_{L^{*}}\leq 1$. By homogeneity, we get $\left<u,v\right>\leq 0$ $\lambda$-a.e. for any bounded $v$, which is only possible when $u=0$ $\lambda$-a.e., similarly to the Luxemburg pseudo-norm.
@@
$=$ $\inf\Big\{\alpha r^{\prime}>0:\int L(\alpha u/r^{\prime})\,d\lambda\leq 1\Big\}\,=\,\alpha\,\|u\|_{L}.$
@@
Integrating both sides and using $\int L(u_{1}/r_{1})\,d\lambda=\int L(u_{2}/r_{2})\,d\lambda=1$, we get
@@ 9. Relationship Between Luxemburg and Orlicz Pseudo-Norms
First let us emphasize the following natural bound on the integrals in (8.2). We keep the same assumptions about the convex function $L$ as in the previous section.
@@
Step 1: Reduction of (9.4) to bounded functions $u$. Consider the sequence $u_{k}$ defined to be $u$ if $|u|\leq k$ and to be zero otherwise. Putting $\Omega_{k}=\{|u|\leq k\}$, we have, by (8.3),
@@
If $\int L^{*}(v)\,d\lambda\leq 1$, that is, when $\|v\|_{L^{*}}\leq 1$, recall that, by the definition (7.2),
@@
In the other case where $c=\int L^{*}(v)\,d\lambda>1$, consider the function $\widetilde{v}=v/c$. Since $L^{*}$ is convex and $L^{*}(0)=0$, we have $L(\alpha x)\leq\alpha L(x)$ for all $x\in{\mathbb{R}}^{n}$ and $0\leq\alpha\leq 1$. In particular,
@@ 10. Perturbations of the Luxemburg Pseudo-Norm
$c)$ For any measurable function $u:\Omega\rightarrow{\mathbb{R}}^{n}$, $\|u\|_{L(\lambda)}\leq\|u\|_{L_{\varepsilon}(\lambda)}$ for all $\varepsilon>0$, and if $u$ is bounded, then $\|u\|_{L_{\varepsilon}(\lambda)}\downarrow\|u\|_{L(\lambda)}$.
$d)$ For any measurable function $u:\Omega\rightarrow{\mathbb{R}}^{n}$, $\|u\|_{L_{\varepsilon}^{*}(\lambda)}\leq\|u\|_{L^{*}(\lambda)}$ and $\|u\|_{L_{\varepsilon}^{*}(\lambda)}\uparrow\|u\|_{L^{*}(\lambda)}$.
@@
For the last claim, put $r=\|u\|_{L^{*}}$. There is nothing to prove, if $r=0$. Assuming that $r$ is positive and finite, we have $\int L^{*}(u/r)\,d\lambda=1$ and $\int L^{*}(u/r^{\prime})\,d\lambda>1$ whenever $0<r^{\prime}<r$. By $b)$ and applying the monotone convergence theorem, it follows that for such values of $r^{\prime}$

### v3

@@ 2. Differentiability and Subgradients of Convex Functions
Hence this set is Borel measurable, and the gradient function $\nabla L(x)$ is Borel measurable as well on $E$. In fact, the function $\nabla L$ is continuous on $E$ (cf. [12]). This implies that if $L$ is differentiable everywhere, then it is $C^{1}$-smooth. For various results about convex functions, let us also refer to the book [3].
@@
$|y|\,=\,\frac{1}{r}\,\sup_{|h|\leq r}\left<y,h\right>\,\leq\,\frac{1}{r}\,\sup_{|h|\leq r}(L(x+h)-L(x)),\quad r>0.$ (2.2)
@@ 3. Existence of Measurable Subgradients
This means that the real-valued function $x\rightarrow h_{T(x)}(\theta)$ is upper semi-continuous on $\Omega$ for any $\theta\in{\mathbb{R}}^{n}$, or equivalently, for any $\theta\in S^{n-1}$. Using the identification of ${\mathbb{K}}_{n}$ and $\widetilde{\mathbb{K}}_{n}$, we may consider a map $\widetilde{T}:\Omega\rightarrow\widetilde{\mathbb{K}}_{n}$ and say that it is upper semi-continuous, if for any $\theta\in S^{n-1}$,
@@
$C=\big\{f\in C(S^{n-1}):f(\theta_{i})<c_{i}\ \ i=1,\dots,n\big\},\quad\theta_{i}\in S^{n-1},\ c_{i}\in{\mathbb{R}}.$
@@ 4. Remarks on the Construction of Measurable Subgradients
As mentioned in [4] with reference to some recent and old publications, evaluating a subgradient may be a challenging task; this difficulty has motivated the development of numerous subdifferential approximations. In Proposition 3.1 one of those constructions is described using formula (3.1). However, a preferable approach would involve directional derivatives $L(x,\theta)$ defined in (2.3) (which are treated as more accessible quantities). The authors of [4] consider and solve this problem in dimension $n=2$.
@@
Hence, by Jensen’s inequality,
@@ 5. The $\Delta_{2}$-Condition
First, let us note that the $\Delta_{2}$-condition (5.1) is equivalent to the property that the function
@@
$c)$ $R(rs)\leq R(r)R(s)$ for all $r\geq 0$, $s\geq 1$;
$d)$ $R$ has a sub-polynomial growth: putting $p=R^{\prime}(1+)$, we have
$R(r)\leq r^{p}\ (r\geq 1).$ (5.4)
Proof. a)b)a)-b) are clear. For $c)$, given $x\in{\mathbb{R}}^{n}$ with $|x|\geq 1$ and $r\geq 0$, $s\geq 1$, we have
$\frac{L(rsx)}{L(x)}=\frac{L(r(sx))}{L(sx)}\frac{L(sx)}{L(x)}\leq R(r)R(s),$
where we used $|sx|\geq 1$. Taking the supremum over all admissible $x$ on the left-hand side, the claim in $c)$ follows.
This property implies that, for any real $a\geq 1$ and any integer $m\geq 0$,
@@
Hence in the limit (5.6) yields (5.4). ∎
Note that the polynomial bound in (5.4) implies
@@
with constant $C=\sup_{|x|=1}L(x)$.
One can now turn to the characterization of the $\Delta_{2}$-condition. Recall that $L^{\prime}(x,x)$ denotes the directional derivative of $L$ at the point $x$ in the direction of $x$, according to (2.3).
@@
In this case, both suprema are at most $p=R_{L}^{\prime}(1+)$. If $L$ is differentiable, this condition simplifies to
@@
Let us now explain why the finiteness ofboth suprema in (5.7) is necessary for the $\Delta_{2}$-condition. Assume that (5.1) holds. By (5.4),
@@
By the convexity of $L$, for all $h\in{\mathbb{R}}^{n}$ and $y\in\partial L(x)$,
@@ 6. A Two-Sided $\Delta_{2}$-Condition
Again, let $L$ be a non-negative convex function on ${\mathbb{R}}^{n}$ such that $L(0)=0$, $L(x)>0$ for $x\neq 0$. In order to describe the possible behavior of $L(x)$ for large $x$, Krasnosel’skii and Rutickii also discussed in [5] related properties such as the $\Delta^{\prime}$-condition and the $\Delta_{3}$-condition. Here we focus on a formally different class of convex functions, whose behaviour can be controlled not only at infinity, but also around zero.
@@
It was shown in [5] that, if $L$ satisfies (6.1) for all $x$ large enough (in dimension $n=1$), it is possible to redefine $L$ near zero in such a way that this inequality would hold for all $x$. The new function is thus equivalent to the original one (in the sense of [5]), and therefore both functions lead to the identical associated Orlicz spaces. Nevertheless, this two-sided $\Delta_{2}$-condition introduces some new remarkable features to the analysis of convex functions.
@@
is finite on the positive half-axis. We call it the Young function associated to $L$. Basic properties of $\Phi_{L}$ are almost identical to those of $R_{L}$.
@@
Properties a)b)a)-b) are obvious. The proof of $c)$ and of the second inequality in (6.3) is similar to the one from Proposition 5.1. As for the first inequality, note that, due to the improved property $c)$, the bound (5.5) holds true for all $a\geq 0$. If $0<r<1$ and $0<a<1$, we apply this bound with $m=[\frac{\log r}{\log a}]$, so that $a^{m}\geq r$. Using the monotonicity of $R$, it gives
$R(r)\,\leq\,R(a)^{m}\,\leq\,\exp\Big\{\Big(\frac{\log r}{\log a}-1\Big)\log R(a)\Big\}\,=\,\frac{1}{R(a)}\,\exp\Big\{\frac{\log R(a)}{\log a}\log r\Big\}.$
For $a=1-\varepsilon$ with $\varepsilon\downarrow 0$, we have
$R(a)=1-R^{\prime}(1-)\,\varepsilon+o(\varepsilon),\quad\log R(a)=-R^{\prime}(1-)\,\varepsilon+o(\varepsilon).$
Therefore, in the limit the above inequality yields the first bound in (6.3).
Now, as a closely related function, also introduce
@@
Let us record without proof its basic properties.
@@
Necessarily $p=p_{-}=p_{+}$. Typical examples of such functions are provided by $L(x)=\|x\|^{p}$, $x\in{\mathbb{R}}^{n}$, with parameter $p\geq 1$, where $\|\cdot\|$ is an arbitrary norm on ${\mathbb{R}}^{n}$.
@@ 7. Characterization of the Two-Sided $\Delta_{2}$-Conditions
Proof. Suppose that $p=p_{-}>1$ and let $r\geq 1$, $x\in{\mathbb{R}}^{n}$. Applying the first inequality in (6.3), i.e. $L(sx)\leq s^{p}L(x)$, $s\in[0,1]$, we get
@@
Changing the variable $rx=y$, we get $L^{*}(r^{p-1}y)\leq r^{p}L^{*}(y)$. The replacement $r^{p-1}=s$ yields $L^{*}(sy)\leq s^{q_{+}}L^{*}(y)$ or equivalently
@@
This implies that $L^{*}$ satisfies the two-sided $\Delta_{2}$-condition and also proves the first inequality in (7.3).
A similar argument based on the second inequality in (6.3) leads to
@@
which gives the second inequality in (7.3). ∎
Remark 7.3. If $L(x)\geq c|x|$ for all $x\in{\mathbb{R}}^{n}$ with some constant $c>0$, then $L^{*}(x)=0$ for all sufficiently small $|x|$. Hence, in this case we may not speak about the $\Delta_{2}$-condition for the Legendre transform of $L$. For an illustration, let us return to Example 6.5 with $p=1$ and consider the convex function
@@ 8. Luxemburg and Orlicz Pseudo-Norms for Vector-Valued Functions
when the value of $r$ under the infimum sign exists, and $\|u\|_{L}=\infty$ otherwise.
@@
Note that $\|u\|_{L}=0$ implies that $u=0$ with $\lambda$-probability 1. Indeed, suppose that $u\neq 0$ on a set $A\subset\Omega$ of positive $\lambda$-probability. Then, using $L(x)>0$ for all $x\neq 0$, we would have $L(\alpha u)\uparrow\infty$ as $\alpha\uparrow\infty$ on the set $A$, so that $\int L(\alpha u)\,d\lambda\uparrow\infty$ or equivalently $\int L(u/r)\,d\lambda\uparrow\infty$ as $r\downarrow 0$. According to (8.1), we then get $\|u\|_{L}>0$.
@@
for any $r>0$, and hence $\|\widetilde{v}\|_{L^{*}}\leq\|v\|_{L^{*}}$. So, after this modification of $v$ the integrals in (8.2) may only increase, while the norm of the modified function $\widetilde{v}$ may only decrease. Therefore,
@@
Note that $|u|_{L}=0$ if and only if $\left<u,v\right>\leq 0$ $\lambda$-a.e. whenever $\|v\|_{L^{*}}\leq 1$. By homogeneity of this pseudo-norm, we get $\left<u,v\right>\leq 0$ $\lambda$-a.e. for any bounded $v$, which is only possible when $u=0$ $\lambda$-a.e., similarly to the Luxemburg pseudo-norm.
@@
$=$ $\inf\Big\{\alpha r^{\prime}>0:\int L(u/r^{\prime})\,d\lambda\leq 1\Big\}\,=\,\alpha\,\|u\|_{L}.$
@@
Integrating both sides over $\lambda$ and using $\int L(u_{1}/r_{1})\,d\lambda=\int L(u_{2}/r_{2})\,d\lambda=1$, we get
@@ 9. Relationship Between Luxemburg and Orlicz Pseudo-Norms
First let us emphasize the following natural bound on the integrals in (8.2). We keep the same assumptions about the convex function $L$ as in the previous section. As is standard, put $a^{+}=\max(a,0)$ for $a\in{\mathbb{R}}$.
@@
Step 1: Reduction of (9.4) to bounded functions $u$. Consider the sequence of functions $u_{k}$ defined to be $u$ if $|u|\leq k$ and to be zero otherwise. Putting $\Omega_{k}=\{|u|\leq k\}$, we have, by (8.3),
@@
If $\int L^{*}(v)\,d\lambda\leq 1$, that is, when $\|v\|_{L^{*}}\leq 1$, then, by the definition (8.2),
@@
In the other case where $c=\int L^{*}(v)\,d\lambda>1$, consider the function $\widetilde{v}=v/c$. Since $L^{*}$ is convex and $L^{*}(0)=0$, we have $L^{*}(\alpha x)\leq\alpha L^{*}(x)$ for all $x\in{\mathbb{R}}^{n}$ and $0\leq\alpha\leq 1$. In particular,
@@ 10. Perturbations of the Luxemburg Pseudo-Norm
$c)$ Given a measurable function $u:\Omega\rightarrow{\mathbb{R}}^{n}$, we have $\|u\|_{L(\lambda)}\leq\|u\|_{L_{\varepsilon}(\lambda)}$ for all $\varepsilon>0$, and if $u$ is bounded, then $\|u\|_{L_{\varepsilon}(\lambda)}\downarrow\|u\|_{L(\lambda)}$.
$d)$ Given a measurable function $u:\Omega\rightarrow{\mathbb{R}}^{n}$, we have
$\|u\|_{L_{\varepsilon}^{*}(\lambda)}\leq\|u\|_{L^{*}(\lambda)}\quad{\sl and}\quad\|u\|_{L_{\varepsilon}^{*}(\lambda)}\uparrow\|u\|_{L^{*}(\lambda)}.$
@@
For the last claim in $d)$, put $r=\|u\|_{L^{*}}$. There is nothing to prove, if $r=0$. Assuming that $r$ is positive and finite, we have $\int L^{*}(u/r)\,d\lambda=1$ and $\int L^{*}(u/r^{\prime})\,d\lambda>1$ whenever $0<r^{\prime}<r$. By $b)$ and applying the monotone convergence theorem, it follows that for such values of $r^{\prime}$